The ages of a mother and daughter add up to 66. The Mother's age is the daughter's age reversed. How old could they be?
Answer by edjones(8007)

m+d=66
d=10t+u
m=10u+t
10t+u+10u+t=66
llu+11t=66
u+t=6
.
24 and 42 or
15 and 51 or
06 and 60 (unlikely)
.
Ed